OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. One side of a triangle is 4 cm shorter than the base, x. The other side is 6 cm longer than the base. What lengths of the base will allow the perimeter of the triangle to be at least 29 cm?

OpenStudy (callisto):

The 3 sides are x, x-4, and x+6 Perimeter = x + (x-4) + (x+6) = 3x +2 Perimeter of the triangle to be at least 29 cm \[ perimeter \ge 29\]\[ 3x +2 \ge 29\] Can you solve it from here?